OpenStudy (anonymous):

calculate the molarity & normality of a solution containing 0.5gm of NAOH dissloved in 500ml solution.

OpenStudy (anonymous):

@DLS ,@sauravshakya @saifoo.khan

OpenStudy (anonymous):

@sauravshakya

OpenStudy (dls):

\[\LARGE Molarity=\frac{Number \of\moles}{Volume}\]

OpenStudy (dls):

Can you calculate the number of moles of NaOH in 0.5 grams?

OpenStudy (australopithecus):

Molarity = \[Molarity = \frac{Moles}{LitresOfSolution}\] Normality is used to express acid/base equivalents, so it is referring to the amount of H+ or OH- that can be released from a single molecule, so it is not looking at the molecule as a whole. This is useful for acid base titrations and other reactions. We use this equation to find Normality \[Normality = Molarity(Equivalents)\] To figure out Normality you simply need to divide by equivalents. EXAMPLES: So just say you have 1M of HCl, it has an equivalency of 1 because 1mole of HCl gives 1mole of H+ so, 1M*1equivilant = 1N So just say you have 2M of H2SO4, it has an equivalency of 2 because 1mole of H2SO4 can give off 2mole of H+ so, 4M*2equivialents = 4N So just say you had 4M of H3PO4, it has an equivalency of 3 because 1mole of H3PO4 can give off 3mole of H+ so, 4M*3equivalents = 12N Hope this helps if you have any questions feel free to ask